OpenStudy (anonymous):

2x^-1 y^4 z^0 ( _______________ ) ^-2 3x^-3 y^2

OpenStudy (anonymous):

\[(\frac{2x^{-1}y^4z^0}{3x^{-3}y^2})^{-2}\]\[(\frac{3x^{-3}y^2}{2x^{-1}y^4z^0})^{2}\]\[(\frac{3x^{-3-(-1)}y^{2-4}}{2\times 1})^{2}\]\[(\frac{3x^{-2}y^{-2}}{2})^{2}\]\[(\frac{3}{2x^{2}y^{2}})^{2}\]\[\frac{3^2}{2^2x^{2*2}y^{2*2}}\]\[\frac{9}{4x^{4}y^{4}}\]

OpenStudy (mertsj):